OpenStudy (anonymous):

hat is the solution to the equation? 8x = 32 A. x = 256 B. x = 3 C. x = 24 D. x = 4

OpenStudy (anonymous):

@amandasoto144

OpenStudy (harsimran_hs4):

divide both sides by 8 , what do you get ?

OpenStudy (anonymous):

0.25

OpenStudy (anonymous):

The answer is D. x=4

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i was not done i that what i get when you divide by 8

OpenStudy (anonymous):

you cant give me medal when i did not do nothin g

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

yes the answer is 8, i divided the opposite

OpenStudy (anonymous):

i mean 4

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Which word phrase represents the variable expression? 3n + 6 A. three more than six times a number B. six more than the sum of a number and three C. three times the sum of a number and six D. six more than three times a number

OpenStudy (anonymous):

i THIN K b

OpenStudy (anonymous):

THINK

OpenStudy (anonymous):

yea its b

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

(4 pt) Which is an equation for the given word problem? After Joel deposits $25.50 into his savings account, his balance is $275.00. Which equation can be used to find the amount, n, in his account before he made the deposit? A. n + 275.00 = 25.50 B. 25.50 + n = 275.00 C. 25.50 – n = 275.00 D. n – 275.00 = 25.50

OpenStudy (anonymous):

last answer was D

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

we know its not A because it would give a negative balance and D would give a even higher balance

OpenStudy (anonymous):

C would also give a negative number so its B because you subtract the deposit put it

OpenStudy (anonymous):

x=4